# The ECME Algorithm Using Factor Analysis for DOA Estimation in Nonuniform Noise

### v2

@@ 5 Numerical Results
Figure 2: Scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from the FAAN method and the ECME algorithm.
Figure 3: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from the FAAN method and the ECME algorithm.
@@ 6 Conclusion
In this article, the ECME algorithm, an extension of the EM algorithm, has been designed for MLFA, which has almost the same computational complexity at each iteration as the FAAN method with excellent convergence properties. However, numerical results have shown that the ECME algorithm yields faster stable convergence and is computationally more efficient. Importantly, the signal subspace estimated by the ECME algorithm can be employed for the subspace based DOA estimation in unknown nonuniform noise.

### v3

@@ 5.2 Accuracy Evaluation
Figure 5: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from both algorithms versus $L$
Figure 6: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from both algorithms versus $\gamma$
Figure 7: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2},\hat{\theta}_{3})$ from both algorithms versus $\gamma$
Figure 8: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from both algorithms versus NN
Fig. 9 and Fig. 10 compare the ECME algorithm with the state-of-the-art subspace separation approach in [Majdoddin], called “Good Method”. In Fig. 9, $\mathbf{P}=\gamma\times\mathbf{I}_{M}$ while in Fig. 10, $M=3$, $\theta_{1}=40^{\circ}$, $\theta_{2}=80^{\circ}$, $\theta_{3}=120^{\circ}$. We can observe that each RMSE curve of DOA estimation from the ECME algorithm is above that from the “Good Method”, so the noise subspace estimated by the “Good Method” is more accurate. Furthermore, because the “Good Method” only requires $5$ iterations and its computational complexity of each iteration also involves $\mathit{\mathcal{O}}(N^{3})$, it is computationally more efficient. As a consequence, MLFA is not the best choice for the subspace based DOA estimation in unknown nonuniform noise [^6] It is worth pointing out that the ECME algorithm is valuable since this algorithm should be the best choice for MLFA in other applications, e.g., psychometrics, econometrics, statistics and so on..
Figure 9: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from the ECME algorithm and a method versus $\gamma$
Figure 10: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2},\hat{\theta}_{3})$ from the ECME algorithm and a method versus $L$
@@ 6 Conclusion
In this article, the ECME algorithm, an extension of the EM algorithm, has been designed again for MLFA by introducing new complete data, which can thus use two explicit formulas to sequentially update the estimates of parameters at each iteration and have excellent convergence properties. Theoretical analysis has shown that the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method. However, numerical results have shown that the ECME algorithm yields faster stable convergence and the convergence to the global optimum is easier. Importantly, MLFA is not the best choice for the subspace based DOA estimation in unknown nonuniform noise.